the hulk kicks a soccer ball off with an initial velocity of 84m/s (30 degrees universe horizontal.
A. What is the balls horizontal velocity?
B. What is the balls flight time?
C. What is the balls maximum height?
D. How far forward does the ball hit the ground?
E. What is the balls velocity when it hits the ground?
1
Expert's answer
2019-06-24T09:23:24-0400

A) The horizontal velocity equals to:


"v_{hor}=v*\cos{30^{\circ}}=84*\cos{30^{\circ}}=73 m/s"


B) The ball flight time:


"t_f=\frac{2v_0*\sin{30^{\circ}}}{g}=\frac{2*84*\sin{30^{\circ}}}{9.8}=8.57 s"


C) The maximum height:


"H=\frac{(v_0*\sin{30^{\circ}})^2}{2g}=90 m"


D) The length of the flight is:


"L=v_{hor}*t_f=73*8.57=625.6 m"


E) According to the energy conservation principle, final velocity equals to the initial velocity:


"v_f=v_0=84 m\/s"